# Semantic Regexes: Auto-Interpreting LLM Features with a Structured Language

### v1

@@ C.1 Feature Description Methods
We evaluate three approaches for generating natural language explanations of LLM features: token-act-pair (Bills et al., 2023), max-acts (Paulo et al., 2024), and our proposed semantic-regex method. All methods use gpt-4o-mini as the explainer model. We follow Neuronpedia’s reference implementations (Lin, 2023) for each baseline method. Implementation details for each method are provided below, and a summary of the settings is shown in Table A2.
@@
We follow Neuronpedia’s implementation (Lin, 2023) of the feature description method proposed in Bills et al. (2023), called oai_token-act-pair in the Neuronpedia interface. The explainer model is prompted with the original feature description system instructions and the three few-shot examples used in their paper (LABEL:lst:oai-prompt). For each feature, we supply the top five activating data examples from Neuronpedia (Lin, 2023), where each example consists of up to a 64-token window centered on the maximally activating token. Tokens are presented in the format:
@@ A Semantic Regex is a structured pattern composed of:
* @{:context C:}(S) - matches S that only activates in the context C (e.g., @{:context political:}([:lexeme run:]) matches "she<< ran>> for office" and "<<running>> for govenor" but not "I<< run>> marathons").
@@
[EXPLANATION]: Nouns representing a distinct objects that contains something, sometimes preciding a quotation mark.
@@
* @{:context C:}(S) - matches S that only activates in the context C (e.g., @{:context political:}([:lexeme run:]) matches "she<< ran>> for office" and "<<running>> for govenor" but not "I<< run>> marathons").
@@
5. Do not includer the delimiters tokens (<< >>) in your Semantic Regex.
@@ 6. Output a short explanation followed by "SR: " and then the Semantic Regex
C.2 Evaluation Metrics
We evaluate the feature description metrics across a suite of evaluation metrics. We take these evluation metrics from Paulo et al. (2024) (Eleuther) and Puri et al. (2025) (FADE). We follow their implementations closely, but make some modifications for consistency across metrics. To account for the semantic regex structure, we make small adjustments to the original prompts and inject a description of the semantic regex language (full prompts in LABEL:lst:detection-prompt-nl, 12, LABEL:lst:detection-prompt-sr, 12, LABEL:lst:fuzzing-prompt-nl, 12, LABEL:lst:fuzzing-prompt-sr, 12, LABEL:lst:fade-prompt-nl, 12, LABEL:lst:fade-prompt-sr, 12, LABEL:lst:clarity-prompt-nl, 12 and LABEL:lst:clarity-prompt-sr). The hyperparameters are listed in Table A3.
All metrics use gpt-4o-mini as the evaluator model. Metrics typically rely on a set of activating examples (positives) and random examples (negatives). Unless otherwise specified, we use 50 activating examples and 50 negative examples per feature, each consisting of a 32-token window centered on the maximally activating token. Activating example sampling varies per metric, but all random examples are sampled from alternative features in Neuronpedia; given the number and diversity of features, this approximates random dataset sampling. To evaluate the description, many of the methods compute the feature’s activation on these data sets. Following prior work (Paulo et al., 2024), when we compute activations, we ignore the beginning-of-sequence tokens.
@@ C.2.1 Eleuther Metrics
Random examples are highlighted using the activation pattern of the positive examples, which ensures an equal distribution of activating tokens across both example sets. The evaluator is asked whether the description matches the highlighted regions, and the score is computed as the balanced accuracy of these judgments relative to the ground-truth labels. This yields a stricter variant of detection by focusing the match decision on the most activating regions.
C.2.2 FADE Metrics
@@
Unlike the other metrics, clarity evaluates a feature description by testing whether it can generate activating text. Here, the evaluator model is prompted with the description and asked to generate candidate examples (full prompts are shown in LABEL:lst:clarity-prompt-nl and LABEL:lst:clarity-prompt-sr). We make 10 independent generation calls, producing a set of positives, and sample an equal number of random examples as negatives. Each example is then passed through the subject model to obtain its activation value. clarity is computed as the Gini index (a rescaled ROC AUC), measuring whether generated examples achieve higher activations than the random negatives.
@@
faithfulness measures how well the feature description captures the feature’s causal role in the LLM. Unlike the other metrics that are based on analyzing activating data, faithfulness compares the feature description against steered model generations where the feature is amplified or ablated. Here, we sample 10 random examples. Then we have the model continue each example by generating 30 additional tokens. We apply this strategy multiple times, where each time the feature is ablated (set to 0) or amplified. In the amplified settings, we set its strength equal to its maximum known data activation multiplied by its modification factor (1, 10, 100). Given the steered generations and the feature description, we follow the same rating task as in responsiveness and purity, asking the evaluator model to provide a discrete judgment (0, 1, or 2) indicating how strongly the text matches the description. Following FADE, samples labeled as 1 are discarded before scoring. The evaluator model prompts are shown in LABEL:lst:fade-prompt-nl and LABEL:lst:fade-prompt-sr. We compute faithfulness as the maximum proportion of matching generations at each modification factor compared to the proportion of matching generations when the feature is ablated.
@@
ou are an intelligent and meticulous linguistics researcher.
@@
Appendix D Cost Analysis
The cost to compute semantic-regex feature descriptions is similar to the cost of prior methods. If we were to compute feature descriptions for every feature in GPT-2-RES-25k, we could compute semantic regexes for \$65.27 using gpt-4o-mini. This is on par with prior methods, where token-act-pair would cost \$67.67 and max-acts would cost \$54.01.
The description cost per feature is determined by the number of input tokens (system prompt and few-shot examples, $T_{\text{prompt}}$, the feature tokens, $T_{\text{feature}}$) and the generated output tokens ($T_{\text{out}}$):
$\text{Cost per Feature}\;=\;P_{\text{in}}(T_{\text{prompt}}\;+\;T_{\text{feature}})\;+\;P_{\text{out}}(T_{\text{out}})$ (1)
where $P_{\text{in}}$ and $P_{\text{out}}$ are the API prices for input and output tokens.
We report the token counts and resulting costs for each feature description method in A4. The relative differences in the number of tokens are the result of token-act-pair using more verbose example formatting than max-acts and semantic-regex, semantic-regex using fewer top activating examples max-acts, and semantic-regex adding the additional semantic regex description to the input prompt. At the time of writing, the OpenAI API costs [^3] https://platform.openai.com/docs/pricing \$0.15/\$0.60 USD per million input/output tokens for gpt-4o-mini and \$2.50/\$10.00 USD per million input/output tokens for gpt-4o. These prices assume no input caching, which can vary due to resource availability, and thus are an upper bound on the total price.
Table A4: Cost analysis for feature description generation is based on the number of input tokens (system prompt, few-shot examples, and feature input) and generated output tokens. We report the mean $\pm$ std number of tokens in the feature input and generated output over 130 features from gpt2-small (10 per layer). Tokenization is computed using OpenAI’s gpt-4o model family tokenizer.
token-act-pair max-acts semantic-regex
[…]

### v2

@@ F.1 Feature Description Methods
We evaluate three approaches for generating descriptions of LLM features: token-act-pair (Bills et al., 2023), max-acts (Paulo et al., 2024), and our proposed semantic-regex method. All methods use gpt-4o-mini as the explainer model. We follow Neuronpedia’s reference implementations (Lin, 2023) for each baseline method. Implementation details for each method are provided below, and a summary of the settings is shown in Table A4.
@@
We follow Neuronpedia’s implementation (Lin, 2023) of the feature description method proposed in Bills et al. (2023), called oai_token-act-pair in the Neuronpedia interface. The explainer model is prompted with the original feature description system instructions and the three few-shot examples used in the original paper (LABEL:lst:oai-prompt). For each feature, we supply the top five activating data examples from Neuronpedia (Lin, 2023), where each example consists of up to a 64-token window centered on the maximally activating token. Tokens are presented in the format:
@@ A Semantic Regex is a structured pattern composed of:
* @{:context C:}(S) - matches S that only activates in the context C (e.g., @{:context political:}([:lexeme run:]) matches "she<< ran>> for office" and "<<running>> for governor" but not "I<< run>> marathons").
@@
[EXPLANATION]: Nouns representing a distinct object that contains something, sometimes preceding a quotation mark.
@@
* @{:context C:}(S) - matches S that only activates in the context C (e.g., @{:context political:}([:lexeme run:]) matches "she<< ran>> for office" and "<<running>> for governor" but not "I<< run>> marathons").
@@
5. Do not include the delimiters tokens (<< >>) in your Semantic Regex.
@@ 6. Output a short explanation followed by "SR: " and then the Semantic Regex
F.2 Evaluation Metrics
We evaluate the feature description metrics across a suite of evaluation metrics. We take these evaluation metrics from Paulo et al. (2024) (Eleuther) and Puri et al. (2025) (FADE). We follow their implementations closely, but make some modifications for consistency across metrics. To account for the semantic regex structure, we make small adjustments to the original prompts and inject a description of the semantic regex language (full prompts in LABEL:lst:detection-prompt-nl, 12, LABEL:lst:detection-prompt-sr, 12, LABEL:lst:fuzzing-prompt-nl, 12, LABEL:lst:fuzzing-prompt-sr, 12, LABEL:lst:fade-prompt-nl, 12, LABEL:lst:fade-prompt-sr, 12, LABEL:lst:clarity-prompt-nl, 12 and LABEL:lst:clarity-prompt-sr). The hyperparameters are listed in Table A5.
All metrics use gpt-4o-mini as the evaluator model. Metrics typically rely on a set of activating examples (positives) and random examples (negatives). Unless otherwise specified, we use 50 activating examples and 50 negative examples per feature, each consisting of a 32-token window centered on the maximally activating token. Activating example sampling varies per metric, but all random examples are sampled from alternative features in Neuronpedia; given the number and diversity of features, this approximates random dataset sampling. To evaluate the description, many of the methods compute the feature’s activation on these data sets. Following prior work (Paulo et al., 2024), we ignore the beginning-of-sequence tokens when computing activations.
@@ F.2.1 Eleuther Metrics
Random examples are highlighted using the activation pattern of the positive examples, which ensures an equal distribution of activating tokens across both example sets. The evaluator is asked whether the description matches the highlighted regions, and the score is computed as the balanced accuracy of these judgments relative to the ground-truth labels. This yields a stricter variant of detection by focusing the match decision on activating regions.
F.2.2 FADE Metrics
@@
Unlike the other metrics, clarity evaluates a feature description by testing whether it can generate activating text. Here, the evaluator model is prompted with the description and asked to generate candidate examples (full prompts are shown in LABEL:lst:clarity-prompt-nl and LABEL:lst:clarity-prompt-sr). We make 10 independent generation calls, producing a set of positives, and sample an equal number of random examples as negatives. Each example is then passed through the subject model to obtain its activation value. clarity is computed as the Gini index, measuring whether generated examples achieve higher activations than the random negatives.
@@
faithfulness measures how well the feature description captures the feature’s causal role in the LLM. Unlike the other metrics that are based on analyzing activating data, faithfulness compares the feature description against steered model generations where the feature is amplified or ablated. Here, we sample 10 random examples. Then we have the model continue each example by generating 30 additional tokens. We apply this strategy multiple times, where each time the feature is ablated (set to 0) or amplified. In the amplified settings, we set its strength equal to its maximum known data activation multiplied by a modification factor (1, 10, 100). Given the steered generations and the feature description, we follow the same rating task as in responsiveness and purity, asking the evaluator model to provide a discrete judgment (0, 1, or 2) indicating how strongly the text matches the description. Following FADE, samples labeled as 1 are discarded before scoring. The evaluator model prompts are shown in LABEL:lst:fade-prompt-nl and LABEL:lst:fade-prompt-sr. We compute faithfulness as the maximum proportion of matching generations at each modification factor compared to the proportion of matching generations when the feature is ablated.
@@
You are an intelligent and meticulous linguistics researcher.